# Spatio-temporal Fourier Transformer (StFT) for Long-term Dynamics Prediction

### v1

@@ 4.3 Main Results
Figure 4: Results of autoregressive prediction in $L_{2}$ relative error across the timespan: (top) perturbed parallel vector potential $\delta A_{\parallel}$ in plasma MHD, and (bottom) magnitude of velocity in shallow-water equation. The shaded region indicates the uncertainty distribution of $3\sigma$ in the relative error of StFT-F.
Table 1: Relative $L_{2}$ error over three spatiotemporal prediction systems, including linear plasma dynamics, Navier-Stokes equations, and shallow-water equations for multi-scale, multi-physics study. AR-StFT refers to the deterministic model’s results. AR-StFT-F denotes the outcomes of the deterministic model with a refinement of flow matching layer for uncertainty quantification, where at each autoregressive step, $50$ sampled are generated to get a mean prediction. The best and second best prediction results for each case are marked in blue and orange.
Dataset Variable(s) AR-StFT AR-StFT-F AR-FNO AR-ViT AR-UNet
Plasma MHD $\delta\phi$ 2.80e-2 2.24e-2 2.28e-1 1.73e-1 1.02e0
$\delta A_{\parallel}$ 2.45e-2 2.30e-2 2.30e-1 3.24e-1 8.13e-1
$\delta B_{\parallel}$ 3.05e-2 2.66e-2 2.33e-1 1.95e-1 7.79e-1
$\delta n_{e}$ 2.84e-2 2.45e-2 2.33e-1 2.08e-1 1.01e0
$\delta n_{i}$ 3.28e-2 2.93e-2 2.33e-1 2.18e-1 1.04e0
$\delta u_{e}$ 3.99e-2 3.73e-2 3.18e-1 2.99e-1 6.96e-1
Navier-Stokes $u$ 3.38e-2 3.30e-2 4.46e-2 5.09e-2 6.16e-2
$v$ 3.60e-2 3.17e-2 4.57e-2 4.60e-2 6.15e-2
$p$ 5.16e-2 4.44e-2 5.90e-2 7.03e-2 7.84e-2
Shallow-Water $\mathbf{V}$ 6.25e-2 6.53e-2 9.53e-2 1.33e-1 2.02e-1
Table 1 shows the forecasting performance of all the models on the three applications. Further details regarding the the problem setup and dataset can be found in Appendix A. We present several test trajectories and visualize their uncertainties across the forecasting time horizon estimated by StFT-F, as illustrated in Figure 6.
@@
StFT performs significantly better than all other baselines across all physical processes in the three applications. In the plasma MHD dataset, the test trajectory has a total of six coupled physics variables and $244$ snapshots. On average, StFT achieves a reduction in error by a factor of six compared to the best baseline, AR-ViT. Although AR-FNO maintains a high resolution in its long-term prediction, it fails to capture the correct dynamics of mode evolution, leading to out-of-phase predictions as shown in Figure 3.
@@
We examine the error growth by plotting the $L_{2}$ relative errors over time, as illustrated in Figure 4 for several representative variables. In the bottom figure for the shallow-water dataset, AR-FNO first appears to slightly better than all other methods during the short term from timestamp 00 to $0.2$, and StFT-F shows superior performance soon after. For the plasma dataset, StFT-F demonstrates dominance starting from $t=0.2$ with a stable performance, while the errors of all baseline methods begin to propagate from that point onward, resulting in a rapid decline compared to StFT-F. Notably, StFT-F exhibits long-term stability relative to the other methods. We also compare the error over time comparing StFT and StFT-F in Figure 5. StFT-F consistently shows improved performance throughout the latter half of temporal domain for the Navier-Stokes dataset. For the plasma dataset, StFT-F begins to prevail from $t=0.6$. These results indicate that our method achieves superior long-term stability and accuracy among all other methods.
For the shallow-water and Navier-Stokes datasets, on average, StFT reduces the errors by an average of $50\%$ and $25\%$, respectively. In Navier-Stokes and plasma datasets, StFT-F not only surpasses StFT but also offers the additional capability of uncertainty quantification, achieving error reductions of $10\%$ in both cases. In the shallow-water dataset, we observe a slight increase in error with StFT-F. More experiment and visualization and results are included in Appendix D.
Figure 5: Results of comparing StFT and StFT-F for the autoregressive prediction in $L_{2}$ relative error across different timestamps. (top) velocity component $u$ in Navier-Stokes equation, and (bottom) magnitude of velocity in shallow-water equation. The shaded region represents the standard deviation distribution of the relative error of StFT-F. In the bottom figure, the uncertainty is negligible that it is not visually discernible. StFT-F demonstrates better performance in the latter stages of the forecasting time horizon compared to StFT.
Figure 6: Evaluation of the forecasting results for three applications: the ground truth, StFT-F prediction, residual of StFT-F prediction, and the associated uncertainty across the time horizon, presented across the initial state (left panel) and final state (right panel) for multiple variables: (a-b) plasma: perturbed electrostatic potential $\phi$ and electron number density $n_{e}$, (c-d) Navier-Stokes: velocity component $u$ and pressure $p$, (e) shallow-water: velocity field $\mathbf{V}$.
@@ 4.4 Uncertainty Quantification
Figure 6 presents the distribution of the empirical standard deviation along with the mean prediction. The generation of test example predictions follows a stochastic recursive process. We generate $S=100$ sample trajectories for each test example, where at each time step $t$, the trajectory sample $u_{t}^{(s)}$ for $s=1,\dots,S$ is computed as: $u_{t}^{(s)}=\mathcal{F}_{\theta_{d}}(\tilde{u}_{t-1}^{(s)})+r_{t}^{(s)}$, where the residual term is sampled from: $r_{t}^{(s)}\sim P_{\theta_{g}}(r|\tilde{u}_{t-1}^{(s)},\mathcal{F}_{\theta_{d}}(\tilde{u}_{t-1}^{(s)}))$. With the generated $S$ trajectory samples, we compute the empirical standard deviation as: $\bar{u}_{t}=\frac{1}{S}\sum_{s=1}^{S}u_{t}^{(s)},$ and $\sigma_{t}=\sqrt{\frac{1}{S}\sum_{s=1}^{S}\left(u_{t}^{(s)}-\bar{u}_{t}\right)^{2}}$. As observed, regions with large errors correspond to those exhibiting significant uncertainties predicted by StFT-F. Additionally, it is evident that uncertainties increase with time. This aligns with our expectation, as errors accumulate during the autoregressive forecasting process. The StFT-F model provides aleatoric uncertainty quantification for studying and analyzing the long-term behaviors of complex physical systems.
4.5 Contribution of Each StFT Block
To assess the contribution of each StFT block corresponding to a specific scale in fitting the training data, we quantify the weight of each block using the following measure:
$\mathcal{W}_{i}=\frac{\|\mathbf{y}_{i}\|_{2}}{\|\mathbf{y}\|_{2}},$ (7)
@@ 4.5 Contribution of Each StFT Block
where $\mathbf{y}_{i}$ represents the prediction from the $i$-th StFT block, and $\mathbf{y}$ denotes the ground truth. We normalize the contributions, and present the contribution of each StFT block for each dataset in Figure 7. A greater contributing factor from the fine-scale layer in StFT is observed in the Navier-Stokes and shallow-water equations, attributed to the sharper changes and smaller scale structures inherent in the dynamics of higher-order nonlinearities. Ablation experiments regarding the impact of multi-scale StFT blocks and frequency path is elaborated in Appendix B.
Figure 7: The contribution of each StFT block across the three datasets. From bottom to top, the patch size decreases. For this plasma data, the first two levels are sufficient to capture the muli-scale structures, whereas for Navier-Stokes and shallow-water equations, the finest scale contributes more significantly.
@@ 5 Conclusion
In this paper, we propose a spatio-temporal Fourier transformer (StFT) for multi-scale and multi-physics long-term dynamics forecasting. Specifically, each StFT block is tailored to address a particular spatial scale, and through a hierarchical composition of multiple StFT blocks spanning different scales, StFT learns the interplay between multiple scales and interactions between multiple physical processes, resulting in stable and accurate long-term dynamics forecasting in an autoregressive manner. Furthermore, we propose and demonstrate the use of a generative residual correction mechanism, which enables meaningful quantification of uncertainties in the predictive model. Our approach demonstrates superior performance compared to several state-of-the-art baselines in the domains of plasma, fluid, and atmospheric dynamics.
@@
Acknowledgement
ZB, DL, SW and LO gratefully acknowledge support from the U.S. Department of Energy, Office of Science, SciDAC/Advanced Scientific Computing Research under Award Number DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract Number DE-AC02-05CH11231.
@@ A.1 Plasma magnetohydrodynamic (MHD) equations
where $u$ and $v$ represent the velocity components in the x and y directions, and $p$ represents the pressure. $f(x,y)$ is the source term, and we set it to $e^{-100((x-0.5)^{2}+(y-0.5)^{2})}$. The Reynolds number is set to 1000. We run a finite difference solver to compute the solutions on a $50\times 50$ spatial grid, with the temporal domain discretized into a total of 101 timestamps over $T\in[0,20]$. We generate a total of 100 trajectories by sampling the four boundary conditions uniformly from $(0.1,0.6)$.
@@ A.3 Spherical shallow-water equations
where $-\pi<\lambda<\pi$, and parameters $\hat{h},\phi_{2},\alpha$, and $\beta$ control the shape and location of the perturbation. The parameters $\alpha$ and $\beta$ are sampled from uniform distributions $\alpha\sim U[0.1,0.5]$ and $\beta\sim U[0.03,0.2]$. We ran the solver from Dedalus (Burns et al., 2020) on a $256\times 256$ spherical grid, and the temporal dimension is discretized into 72 timestamps. We collect a total of 200 trajectories by sampling $\alpha$ and $\beta$.
@@ Appendix B Ablation Study
Table 2 shows the $L_{2}$ relative errors averaged over all the variables. Note that “mono-scale” denotes a single scale of StFT, spanning from the coarsest to the finest level, and “multi-scale” represents the multiple layer setting in the hierarchical structure without frequency path. We find that “multi $+\mathcal{F}$” outperforms both ablation cases. These results demonstrate the effectiveness of the hierarchical composition of StFT blocks and the frequency path in StFT block. With the hierarchical composition, for Plasma MHD, the error drops to $0.0307$ from $0.0805$, and for shallow-water equations (SWE), the error drops to $0.0625$ from $0.0975$. We observe that the fine-level layer setting in the SWE achieves the best performance among the single layer results, and the multi-layer settings further decrease the prediction error. The frequency path in StFT also plays a crucial role, where in Plasma MHD, the error drops to $0.0307$ from $0.0404$, and in the SWE dataset, the error drops to $0.0625$ from $0.0956$.
MHD SWE
mono-scale $+$ $\mathcal{F}$ $0.0805/0.105$ $2.5729/0.101/0.0975$
multi-scale $0.0404$ $0.0956$
multi-scale $+$ $\mathcal{F}$ $\mathbf{0.0307}$ $\mathbf{0.0625}$
@@
Table 2: Ablation study results. We run models with the combinations of StFT blocks and the frequency path. “Mono-scale” denotes a single scale of StFT block from the coarsest to the finest level, and “multi-scale” represents the multiple layer setting in the hierarchical structure. $\mathcal{F}$ stands for the frequency path in StFT blocks.
Appendix C Experimental Details

### v2

@@ 4.3 Main Results
Refer to captionRefer to caption
Figure 4: Left top: Comparing StFT and StFT-F in $L_{2}$ relative error across timestamps for two plasma MHD variables with uncertainty bands of StFT-F. Left bottom: The contribution of each StFT block. From bottom to top, the patch size decreases. For this plasma data, the first two levels are sufficient to capture the multi-scale structures, whereas for Navier-Stokes and shallow-water equations, the finest scale contributes more significantly. Right: Spatiotemporal evolution of perturbed electrostatic potential $\delta\phi$ predicted by autoregressive models. StFT and StFT-F remain accurate past $T=59$, whereas baseline methods exhibit out-of-phase predictions.
@@
Table 1 shows the forecasting performance of all the models on the three applications. We present several test trajectories and visualize their uncertainties across the forecasting time horizon estimated by StFT-F, as illustrated in Figure 5. StFT performs significantly better than all other baselines across all physical processes in the three applications. In the plasma MHD dataset, the test trajectory has a total of six coupled physics variables and $244$ snapshots. On average, StFT achieves a reduction in error by a factor of three compared to the best baseline, AR-DPOT(AFNO). Although AR-FNO maintains a high resolution in its long-term prediction, it fails to capture the correct dynamics of mode evolution, leading to out-of-phase predictions as shown in Figure 4.
@@
We examine the error growth by plotting the $L_{2}$ relative errors over time, as illustrated in Figure 3 for several representative variables. In the bottom figure for the shallow-water dataset, AR-FNO first appears to slightly better than all other methods during the short term from timestamp 0 to $0.2$, and StFT-F shows superior performance soon after. For the plasma dataset, StFT-F demonstrates dominance starting from $t=0.2$ with a stable performance, while the errors of all baseline methods begin to propagate from that point onward, resulting in a rapid decline compared to StFT-F. Notably, StFT-F exhibits long-term stability relative to the other methods. As shown in Figure 4, both of our methods accurately capture the dominant mode phase, and StFT-F generates predictions that are more closely aligned with the ground truth compared to StFT. We also compare the error over time comparing StFT and StFT-F. StFT-F begins to prevail from $t=0.6$. For the shallow-water and Navier-Stokes datasets, on average, StFT reduces the errors by an average of $27\%$ and $25\%$, respectively. In Navier-Stokes and plasma datasets, StFT-F not only surpasses StFT but also offers the additional capability of uncertainty quantification, achieving error reductions of $10\%$ in both cases. In the shallow-water dataset, we observe a slight increase in error with StFT-F. These results indicate that our method achieves superior long-term stability and accuracy among all other methods. More visualizations and results are included in Appendix F.
@@ 4.4 Uncertainty Quantification
Table 2: Average confidence interval (90% and 95%) coverage.
Dataset CI: 90% CI: 95%
Plasma MHD 89.4% 92.5%
Shallow-Water 89.5% 98.0%
Figure 5 presents the distribution of the empirical standard deviation along with the mean prediction. As observed, regions with large errors correspond to those exhibiting significant uncertainties predicted by StFT-F. Additionally, it is evident that uncertainties increase with time. This aligns with our expectation, as errors accumulate during the autoregressive forecasting process. Besides the empirical evaluation relying on the standard deviation plots, we further measure the robustness of StFT’s uncertainty quantification using confidence intervals as shown in Table  2. Specifically, we use the predicted uncertainty to compute empirical coverage by measuring the proportion of ground truth values that fall within a certain confidence interval around the predicted mean. This provides a more rigorous evaluation of StFT’s ability to capture uncertainties. Details about CI coverage and results regarding each variable are included in Appendix  D. For the $90\%$ confidence interval, average coverage is very close to ideal value of $90\%$, demonstrating that StFT-F is well-calibrated around the $90\%$ confidence interval.
Figure 5: Evaluation of forecasting for three applications: ground truth, StFT-F prediction, residual, and uncertainty over time — shown for initial (left) and final (right) states. Variables include: (a) $\delta\phi$ in plasma MHD, (b) $u$ in Navier-Stokes, and (c) $\mathbf{V}$ in shallow-water equations.
4.5 Ablation Study
Table 3: Effect of multi-scale structures and frequency path $\mathcal{F}$.
Model Plasma MHD Shallow-Water
Mono-scale + $\mathcal{F}$ 0.0805 / 0.105 2.5729 / 0.101 / 0.0975
Multi-scale 0.0404 0.0956
Multi-scale + $\mathcal{F}$ 0.0307 0.0625
Multi-scale and frequency path. To evaluate the effectiveness of the hierarchical structure and the frequency path in StFT, we conduct an ablation study employing mono-scale versus multi-scale models, and employing models with and without the frequency path. Experimental results are shown in Table  3.
Table 4: L2 relative error vs. number of scales (plasma MHD).
# Scales L2 Rel. Error
1 0.0805
2 0.0307
3 0.0385
4 0.0391
Convergence study on number of scales. We evaluate the performance of StFT on the number of scales for the plasma MHD dataset. As shown in Table 4, it is observed that increasing the number of scales leads to a reduction in prediction error to a certain point, with the two-layer configuration achieving the lowest error. The inclusion of additional scales results in a slight error increase, indicating a convergence in model performance. Additionally, we conduct an ablation study on the overlapping tokenizer. More details are included in Appendix  B.
4.6 Additional Results
Computational and Space Complexity. We compare StFT with other baselines regarding the inference FLOPS per sample, training time and inference time per batch, and peak memory usage during training. We also compare StFT and StFT without the overlapping tokenizer, and we observed this enhancement comes with negligible impact on compute cost, as in Appendix  C.
@@ 4.6 Additional Results
Contribution of Each Scale. To assess the contribution of each StFT block for a specific scale in fitting the training data, we quantify the weight of each block by: $\mathcal{W}_{i}=\frac{\|\mathbf{y}_{i}\|_{2}}{\|\mathbf{y}\|_{2}}$, where $\mathbf{y}_{i}$ represents the prediction from the $i$-th StFT block, and $\mathbf{y}$ denotes the ground truth. We normalize the contributions, and present the contributions in Figure 4. A greater contributing factor from the fine-scale layer in StFT is observed in the Navier-Stokes and shallow-water equations, attributed to the sharper changes and smaller scale structures inherent in the dynamics of higher-order nonlinearities.
@@ 5 Conclusion
In this paper, we propose a spatio-temporal Fourier transformer (StFT) for multi-scale and multi-physics long-term dynamics forecasting. Specifically, each StFT block is tailored to address a particular spatial scale, and through a hierarchical composition of multiple StFT blocks spanning different scales, StFT learns the interplay between multiple scales and interactions between multiple physical processes, resulting in stable and accurate long-term dynamics forecasting in an autoregressive manner. Furthermore, we propose and demonstrate the use of a generative residual correction mechanism, which enables meaningful quantification of uncertainties in the predictive model. Despite demonstrating superior forecasting ability in SciML, the model is based on regular grids, which constraints its applicability to irregular geometries. As part of future work, we plan to extend the framework to handle irregular domains, broadening its utility for more complex real-world scientific and engineering scenarios. The potential in improving its performance includes model parallelism across the multi-scale StFT blocks and extending StFT-F to end-to-end training.
@@
Acknowledgements
ZB, DL, SW, and LO acknowledge support from the U.S. Department of Energy, Office of Science, SciDAC/Advanced Scientific Computing Research under Award Number DE-AC02-05CH11231 with DL receiving additional support from the Margolis Foundation and MURI AFOSR under grant FA9550-20-1-035. SZ was supported by MURI AFOSR grant FA9550-20-1-035, Margolis Foundation, and NSF Career Award IIS-2046295. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract Number DE-AC02-05CH11231, as well as resources from the NCSA Delta GPU cluster from NCF Access program under award number IIS-2046295.
@@ A.1 Plasma magnetohydrodynamic (MHD) equations
where $u$ and $v$ represent the velocity components in the x and y directions, and $p$ represents the pressure. $f(x,y)$ is the source term, and we set it to $e^{-100((x-0.5)^{2}+(y-0.5)^{2})}$. The Reynolds number is set to 1000. We run a finite difference solver to compute the solutions on a $50\times 50$ spatial grid, with the temporal domain discretized into a total of 101 timestamps over $T\in[0,20]$. We generated a total of 100 trajectories by sampling the four boundary conditions uniformly from $(0.1,0.6)$.
@@ A.3 Spherical shallow-water equations
where $\pi<\lambda<\pi$, and parameters $\hat{h},\phi_{2},\alpha$, and $\beta$ control the shape and location of the perturbation. The parameters $\alpha$ and $\beta$ are sampled from uniform distributions $\alpha\sim U[0.1,0.5]$ and $\beta\sim U[0.03,0.2]$. We ran the solver from Dedalus [6] on a $256\times 256$ spherical grid, and the temporal dimension is discretized into 72 timestamps. We have a total of 200 trajectories by sampling $\alpha$ and $\beta$.
@@
Figure 9: Additional evaluation of forecasting: ground truth, StFT-F prediction, residual, and uncertainty over time - shown for initial (left) and final (right) states. Variables include: (a) $\delta n_{e}$ in plasma MHD, (b) $p$ in Navier-Stokes.
@@ B.1 The Hierarchical Structure and The Frequency Path
Table 5: Ablation study results. We run models with combinations of StFT blocks and the frequency path. $l_{1}$ is the coarsest level, and $l_{3}$ is the finest level. F stands for the frequency path in StFT blocks.
Dataset Setting(s) AR-StFT
Plasma MHD $l_{1}+\mathcal{F}$ 0.0805
$l_{2}+\mathcal{F}$ 0.105
$l_{1}+l_{2}$ 0.0404
$l_{1}+l_{2}+\mathcal{F}$ 0.0307
Shallow-Water $l_{1}+\mathcal{F}$ 2.5729
$l_{2}+\mathcal{F}$ 0.101
$l_{3}+\mathcal{F}$ 0.0975
$l_{1}+l_{2}+l_{3}$ 0.0956
$l_{1}+l_{2}+l_{3}+\mathcal{F}$ 0.0625
Table 5 shows the $L_{2}$ relative errors averaged over all the variables. Note that $l_{1}$ is the coarsest level, and $l_{3}$ is the finest level. $\mathcal{F}$ stands for the frequency path in StFT blocks. We find that multi-scale $\mathcal{F}$” outperforms both ablation cases. These results demonstrate the effectiveness of the hierarchical composition of StFT blocks and the frequency path in StFT block. With the hierarchical composition, for Plasma MHD, the error drops to $0.0307$ from $0.0805$, and for shallow-water equations (SWE), the error drops to $0.0625$ from $0.0975$. We observe that the fine-level layer setting in the SWE achieves the best performance among the single layer results, and the multi-layer settings further decrease the prediction error. The frequency path in StFT also plays a crucial role, where in Plasma MHD, the error drops to $0.0307$ from $0.0404$, and in the SWE dataset, the error drops to $0.0625$ from $0.0956$.
B.2 The Overlapping Tokenizer
We conduct an ablation study on the overlapping tokenizer design, evaluating both predictive performance and computational cost, as reported in Tables 6 and 7.We evaluate the computational cost including the inference FLOPS per sample, training/inference time per batch and peak memory usage. We used a fixed batch size of 20 for all models. The results are summarized in the following two tables. StFT-O stands for StFT with the overlapping tokenizer and StFT-NO for the opposite. All experiments were performed on an NVIDIA A100 GPU. Incorporating the overlapping tokenizer leads to substantial improvements in accuracy, in plasma MHD and shallow-water datasets, with errors reductions of $68\%$ and $10\%$ respectively. Notably, this enhancement comes with negligible impact on computational complexity, inference time and peak memory usage.
Table 6: Effect of the overlapping tokenizer on the prediction error of plasma MHD and shallow-water datasets.
Method Plasma MHD Shallow-Water
With overlapping tokenizer 0.0307 0.0625
Without overlapping tokenizer 0.0986 0.0700
Table 7: Computational complexity comparison of StFT with and without overlapping tokenizer.
Dataset Method GFLOPs Training / Iter (s) Inference / Iter (s) Peak Memory (GB)
MHD StFT-O 0.704 0.181 0.0455 9.49
MHD StFT-NO 0.704 0.157 0.0454 9.49
SWE StFT-O 4.30 0.189 0.0319 12.2
SWE StFT-NO 4.30 0.189 0.0319 12.2
B.3 Convergence on the Number of Scales
We evaluate the performance of StFT on the number of scales for the plasma MHD dataset. Specifically, we run StFT with one scale (patch size of $128$), three scales (patch sizes of 128, 64, and 48), and four scales (patch sizes of 128, 64, 48, and 32). As shown in Table 8, it is observed that increasing the number of scales leads to a reduction in prediction error to a certain point, with the two-layer configuration achieving the lowest error ($0.0307$). Beyond this, the inclusion of additional scales results in a slight increase in error, indicating a convergence in model performance, and further increasing the number of scales may not provide performance gain.
Table 8: Effect of the number of scales on L2 relative error of the overall prediction on the plasma MHD.
Number of Scales L2 Relative Error
1 0.0805
2 0.0307
3 0.0385
4 0.0391
Appendix C Computational and Space Complexity
Figure 10: Comparison of models in terms of mean L2 relative error versus inference time (left) and FLOPs (right) for plasma MHD (top) and shallow-water (bottom).
We compare StFT with other baselines regarding the inference FLOPS per sample, training time per batch, inference time per batch, and peak memory usage (during training). In plasma MHD and shallow-water datasets, compared to FNO, StFT is 2020x smaller, and 88x smaller in computation (FLOPS) respectively. While achieving the highest prediction accuracy, StFT has the same order of magnitude of FLOPs as UNet and ViT in plasma MHD. In the shallow-water equation (SWE), however UNet is much more computationally expensive - 1111x more than StFT – while its relative error is 33x larger than StFT. StFT is also efficient in both training and inference time. For plasma MHD, StFT achieves a $30\%$ reduction in training time and $50\%$ reduction in inference time compared to FNO. For SWE, StFT is $40\%$ faster in training and $60\%$ faster than FNO in inference. Although StFT incorporates dual paths operating in the frequency domain and the spatio-temporal domain, its peak memory usage remains comparable to that of FNO in the plasma MHD, and is reduced by $56\%$ compared to FNO in SWE, and $26\%$ less than UNet in SWE. Figure 10 compares the inference time and FLOPs versus mean L2 relative error for all models.
Table 9: Comparison of methods on plasma MHD and shallow-water datasets regarding the computational and space complexity.
Dataset Method GFLOPs Training / Iter (s) Inference / Iter (s) Peak Memory (GB)
MHD StFT 0.704 0.181 0.0455 9.49
MHD DPOT 1.10 0.0676 0.0120 3.20
MHD FNO 13.8 0.262 0.0900 10.1
MHD UNET 0.462 0.0263 0.00230 1.38
MHD ViT 0.338 0.0171 0.00237 1.65
SWE StFT 4.30 0.189 0.0319 12.2
SWE DPOT 1.61 0.0427 0.0133 8.43
SWE FNO 34.6 0.325 0.0886 30.2
SWE UNET 48.5 0.107 0.0257 16.5
SWE ViT 0.90 0.0130 0.00347 7.45
Appendix D Signficance of UQ on StFT-F
For the Plasma MHD dataset, we measure the average coverage for the first $20$ autoregressively predicted snapshots (each row represents one physical variable), as shown in Table 10. In the last row, we report the average coverage across all physical variables. For the $90\%$ confidence interval, average coverage is very close to the ideal value of $90\%$, demonstrating that StFT-F is well-calibrated around the $90\%$ confidence interval. For the $95\%$ confidence interval, the average coverage is under-confident by $2.5\%$, suggesting the intervals may be slightly narrow. For the shallow-water dataset, we provide the results in Table 11, where we measure the average coverage for the first $10$ predicted snapshots. For the $90\%$ confidence interval, average coverage is also very close to the ideal value of $90\%$, which shows that StFT-F is also well-calibrated for the shallow-water dataset.
Table 10: Confidence interval coverage on the plasma MHD dataset.
Variable CI: 90% CI: 95%
$\delta\phi$ 0.906 0.937
$\delta A_{\parallel}$ 0.977 0.988
$\delta B_{\parallel}$ 0.900 0.930
$\delta n_{e}$ 0.933 0.962
$\delta n_{i}$ 0.867 0.910
$\delta u_{e}$ 0.781 0.823
Average Coverage 89.4% 92.5%
Table 11: Confidence interval coverage on the shallow-water dataset.
Variable CI: 90% CI: 95%
$\mathbf{V}$ 0.895 0.980
Table 12: Training/validation/test data splits and the training budget for all models measured on an A100 GPU.
Dataset Total Split (Train / Val / Test) Training Budget
Plasma MHD 1 traj.
(1,221 snapshots)
927 / 50 / 244 24h
Navier-Stokes 100 traj.
(101 snaps/traj)
80 / 10 / 10 24h
Shallow-water 200 traj.
(72 snaps/traj)
160 / 20 / 20 48h
Appendix E Experimental Details